[figure: 9.7 cm]

Phase diagrams and thermodynamic properties of the Penson–Kolb–Hubbard model: s-wave pairing vs η-paring.

S. Robaszkiewicz and W. R. Czart Institute of Physics, Adam Mickiewicz University Umultowska 85, 61-614 Poznań, Poland

Thermodynamic and electromagnetic characteristics of the Hubbard model with pair– hopping interaction J, i.e. the Penson–Kolb–Hubbard model, are studied for d – dimensional hypercubic lattices (d = 2 (SQ) and d = 3 (SC)). In the analysis we focus on the properties of the superconducting states with Cooper-pair center-of-mass momentum q = 0 (s-phase) and q = Q⃗ (η-phase). The evolutions of the critical temperatures, the gap to T_{c} ratios and the phase diagrams with interaction parameters and band filling are determined. The s-phase is found to exhibit a smooth crossover from the BCS-like limit to the tightly bound pairs regime with increasing pairing strength. On the contrary, the transition into the η-phase correspond to an abrupt change in the ground state structure. For J ≠ 0 both these phases can survive also for repulsive on-site density interaction U (U < U_{c}). For the s-phase and d = 2 SQ lattice the effects of phase fluctuations on the transition temperature are examined within the Kosterlitz– Thouless (KT) scenario. The KT critical temperatures T_{c} and the London penetration depth λ are calculated and the T_{c} vs λ^{−2}(0) plots are derived. Except for weak coupling limit these plots have a shape similar to the experimental Uemura’s plots obtained for various classes of short–coherence length superconductors.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : W. R. Czart

Address for correspondence : Institute of Physics, Adam Mickiewicz University Umultowska 85, 61-614 Poznań, Poland

Email address : czart@amu.edu.pl